If a person is twice as old as his brother. 3 years ago the sum of their ages was 27. Find the age of the brothers.
Answer by algebrahouse.com(1659)

"If a person is twice as old as his brother. 3 years ago the sum of their ages was 27. Find the age of the brothers."

x = one brother
2x = other brother {twice as old as the other one}

x - 3 = one brother 3 years ago
2x - 3 = other brother 3 years ago

x - 3 + 2x - 3 = 27 {sum of ages 3 years ago was 27}
3x - 6 = 27 {combined like terms}
3x = 33 {added 6 to both sides}
x = 11 {divided both sides by 3}
2x = 22 {substituted 11, in for x, into 2x}

Their ages are 11 and 22

For more help from me, visit: www.algebrahouse.com